# Localized Triplons and Site Stuffing in the Quantum Dimer Magnet BiYbGeO5

### v1

@@ V Conclusions
In this work, polycrystalline samples of BiYbGeO5 were synthesized via conventional solid-state reactions, and their magnetic and structural properties were investigated through neutron powder diffraction and inelastic neutron scattering. Neutron diffraction measurements carried out at 20 K, 0.9 K, and 58 mK, confirming the absence of magnetic orderings. Structural refinements revealed significant site disorder, with 19.2(26)% of Yb3+ sites replaced by non-magnetic Bi3+ ions. Inelastic neutron scattering measurements uncovered dispersionless triplon excitations, consistent with localized triplet modes due to isolated Yb–Yb dimers. Gaussian fits to the energy spectrum yielded an XXZ-type anisotropic exchange interaction, with meV and meV. Analysis of the magnetic structure factor further determined the intradimer Yb-Yb distance to be 3.499(11) Å, in excellent agreement with the nearest-neighbor spacing extracted from diffraction. These results confirm that despite substantial structural disorder, BiYbGeO5 hosts a magnetically disordered quantum ground state composed of well-isolated Yb dimers, making it a compelling platform for exploring disorder-resilient quantum magnetism.

### v2

@@ V Conclusions
In this work, polycrystalline samples of BiYbGeO5 were synthesized via conventional solid-state reactions, and their magnetic and structural properties were investigated through neutron powder diffraction and inelastic neutron scattering. Neutron diffraction measurements were carried out at 20 K, 0.9 K, and 58 mK, confirming the absence of magnetic orderings. Structural refinements revealed significant site disorder, with 19.2(26)% of Yb3+ sites replaced by non-magnetic Bi3+ ions. Inelastic neutron scattering measurements uncovered dispersionless triplon excitations, consistent with localized triplet modes due to isolated Yb–Yb dimers. Gaussian fits to the energy spectrum yielded an XXZ-type anisotropic exchange interaction, with meV and meV. Analysis of the magnetic structure factor further determined the Yb-Yb dimer distance to be 3.541(11) Å. These results confirm that despite substantial structural disorder, BiYbGeO5 hosts a magnetically disordered quantum ground state composed of well-isolated Yb dimers, making it a compelling platform for exploring disorder-resilient quantum magnetism.